Risa is 5 years older than Peter. Last year, she was twice as old as Peter was. How old is each now ?
Answer by mananth(16946)

Risa is 5 years older than Peter. Last year, she was twice as old as Peter was. How old is each now ?
...
Peter = x years
Risa = x+5
..
1 year ago
Peter = x-1
Risa (x+5)-1= x+4
..
x+4=2(x-1)
x+4=2x-2
2x-x=4+2
x= 6 Peter's age
Risa = 5+6 = 11 years
...
m.ananth@hotmail.ca